you invested $25,000 in two accounts paying 7% and 9% annual interest,respectively. If the total interest earned for the year was $1770 how much was invested at each rate?
Answer by richwmiller(17219)

Total amount of money invested: $25000
x+y=25000,
Total yearly interest for the two accounts is: $1770
0.07*x+0.09*y=1770
x=25000-y
Substitute for x
0.07*(25000-y)+0.09*y=1770
Multiply out
1750-0.07y+0.09*y=1770
Combine like terms.
0.02*y=20
Isolate y
y=$ 1000.00 at 9%
x=25000-y
Calculate x
x=$24000.00 at 7%
Check
0.07*24000+0.09*1000=1770
1680+90=1770
1770=1770
If this statement is TRUE and neither x nor y is negative then all is well